OpenStudy (anonymous):

3x-5/x^2+4x+4=?/(x+2)^2(x+3) Include steps if possible thankyou(:

hero (hero):

Are you sure you posted this correctly? Something about this doesn't look right

OpenStudy (anonymous):

the questions mark is what your looking for.!

hero (hero):

I guess tutordiffeq can do this one.

OpenStudy (anonymous):

(3x-5)/(x+2)^2= ?/ [(x+2)^2(x+3)] (3x-5)(x+3)= ?

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

x^2+4x+4= (x+2)^2

hero (hero):

I knew that much

hero (hero):

The question mark threw me off

OpenStudy (anonymous):

so then to make the common denominator all you have to do is mulitply the left side by (x+3)/(x+3)

OpenStudy (anonymous):

u good?

hero (hero):

You should have used the drawing board to show your steps

OpenStudy (anonymous):

Thank you.